OpenStudy (anonymous):

x+5y=5 Graph the linear equation and solve for y

OpenStudy (anonymous):

x+5y=5 -x -x --------- 5y=-x+5

OpenStudy (anonymous):

you have to put it in y=mx+b format

OpenStudy (anonymous):

then you simplify

OpenStudy (anonymous):